INAF LBC+MODS queue observing, 2012 May 12-13

Observer: K. Boutsia, M. Fumana
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: D. Thompson
Instrument: MODS + LBC

Summary

Clear skies at the beginning of the night.
~50min lost because of error positioning the mask with MODS. We had to Call R. Pogge to solve this.
20min lost because the rotator hit limits, T.O. cancelled preset and we had to repeat acquisition.
Although the moon was up for the last 2hrs of the night, the sky was a lot darker.
Seeing plot at the end of the log. MODS calibrations finished tonight.

Data Summary

Project_RUN Objects Filter ExpTime Notes
NONINO_12B RXJ1532 dual_grating 5hr15min seeing <1.0" - COMPLETED
NONINO_12D Abell2261 dual_grating 1hr10min seeing ~0.8" - COMPLETED
Standard Feige34 red/dual grating    

Details

11:25 bias script for MODS:
  • ./doBias mods1b 0001-0020 / mods1r 0001-0020

02:54 acqMODS Feige34.acq mods1r 0021-0022

02:59 execMODS Feige34.obs mods1b 0021 / mods1r 0023-0024
  • dimm seeing 0.8"

03:11 acqMODS rxj1532_mods.acq
  • error with the position of the mask
** ERROR: MSELECT MSELECT=13 Move Fault, position at end of move 13 but requested position 14.000000
** Abort, Retry, or Ignore? > Retry
  • reset mselect from command line to recover

03:20 Call Rick for slitmask error - we cannot position the slitmask requested by the project

04:02 acqMODS rxj1532_mods.acq mods1r 0025-0027
  • there is scattered light in the acquisition image
  • this should not affect the science slits, since the light is on the left side and all science slits are mostly on the right

04:16 execMODS rxj1532_mods.obs (ID523416) mods1b 0022-0025 / mods1r 0028-0031

07:21 abort exposure 200sec before the end because of rotator limits
  • we need to wait some minutes for the target to cross the meridian before we start the acquisition again
  • this last exposure is considered completed

07:30 acqMODS rxj1532_mods.acq mods1r 0032-0034

07:42 execMODS rxj1532_mods.obs (ID523416) mods1b 0026-0028 / mods1r 0035-0037

10:04 acqMODS abell2261_mods.acq mods1r 0038-0040

10:17 execMODS abell2261_mods.obs (ID562750) mods1b 0029-0030 / mods1r 0041-0042
  • the second exposure is only 1500sec long, because it was partly done in twilight.
  • Sky was darker today than yesterday and seeing was better so it could be useful

11:31 Abort the second exposure because sky is getting brighter.

11:34 Close Dome and put MODS calibration unit in
  • execMODS abell2261_flats.cal (flats ID562750) mods1b 0033-0037 / mods1r 0045-0049
  • execMODS rxj1532_flats.cal (flats ID523416) mods1b 0038-0042 / mods1r 0050-0054
  • execMODS ~MODSeng/modsScripts/Support/modsSleep.pro

12:16 End of Night

  • Seeing reported by GCS during MODS observations:
    20120513.jpg

-- VincenzoTesta - 12 May 2012
Topic revision: r6 - 14 May 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback